The simples form of (f-g)(x) , given that f(x)= 1-x/2+x, g(x)= 2x-3/x-2 is ?
Answer by stanbon(75887)


(f-g)(x) , given that f(x)= 1-x/2+x, g(x)= 2x-3/x-2 is ?
------------------------------------------
(f-g)(x) = f(x) - g(x)
= (1-x)/(2+x) - (2x-3)/(x-2)
---
Rearrange:
= -(x-1)/(x+2) - (2x-3)/(x-2)
= [-(x-1)(x-2) - (2x-3)(x+2)] / [x^2-4]
= [-(x^2-3x+2) -(2x^2+x-6)] / [x^2-4]
= [-3x^2 + 2x + 4] / [x^2-4]
==================================
Cheers,
Stan H.